OpenStudy (anonymous):

Write the equation of the line that passes through (4, 2) and is parallel to the line y = 2x – 1.

OpenStudy (haseeb96):

first take the slope of the given equation here the slope of given equation is m=2 because in this type of equation always the constant of x is yur slope now it is clear

OpenStudy (anonymous):

but what's the answer?

OpenStudy (haseeb96):

lines are parallel so slope will be same as before m=2

OpenStudy (haseeb96):

now put the point in this equation (y2-y1) = m( x2 -x1) y-2 = 2 ( x - 4) y-2 = 2x -8 2x -y -8 +2 =0 2x - y -6=0

OpenStudy (haseeb96):

this is the answer